Tawnie:

what is 6x divided by 3y

bri05:

what do u think it is?

KingToastyMike:

type out the equation

ItzLily:

6x/3y?

ItzLily:

What is you academic guess and why?

bri05:

i have a math app for this to

cupcakesonice:

its 3

bri05:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cupcakesonice its 3 \(\color{#0cbb34}{\text{End of Quote}}\) not right

cupcakesonice:

yes it is

Tawnie:

idk

cupcakesonice:

its 2

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cupcakesonice its 3 \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @cupcakesonice yes it is \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @cupcakesonice its 2 \(\color{#0cbb34}{\text{End of Quote}}\) Wrong.

bri05:

there are two differnt letters

Tawnie:

im stupid

ItzLily:

Your not stupid..

cupcakesonice:

hey not stupid

Tawnie:

r u sure

supie:

\(6x÷3y\) could also be written as \(\frac{6x}{3y}\) or \(\frac{6x}{3}y\) so the first step would be to divide 6 and 3 what is 6 ÷ 3 ?

ItzLily:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Tawnie im stupid \(\color{#0cbb34}{\text{End of Quote}}\) You just didn't understand it.. your not stupid

bri05:

its 2x/y thats the answer

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bri05 its 2x/y thats the answer \(\color{#0cbb34}{\text{End of Quote}}\) No-

Tawnie:

6/3 is 2

supie:

yes so do you know what the answer would be?

Tawnie:

i think so

supie:

What did u get for the answer

ItzLily:

\[\frac{ 2x }{ y}\] Right

ItzLily:

?

supie:

Close, but not quite.

Tawnie:

yea

Tawnie:

oh

ItzLily:

Can you explain to me?

supie:

All you're doing is dividing the numbers 6 and 2 6x ÷ 3x = \(\frac{6x}{3y}\) which can also be written as \(\frac{6x}{3}y\) all that needs to be divided is 6 and three which equals 2 so you would be left with 2xy

supie:

do you guys understand orrrr-

ItzLily:

Ohh yes I understand!

Tawnie:

ohhh

ItzLily:

I almost had it right.

ItzLily:

Thank you @supie

Tawnie:

i understand now thanks

supie:

np

Tawnie:

:)

M00n123:

6*​x/​(3*​y)

supie:

What?

M00n123:

Hm

Tawnie:

umm

supie:

are you saying that's the answer? if you are then you're wrong

Tawnie:

bye

supie:

bye bye c: